Presented below are the ending balances of accounts for the Kansas Instruments Corporation at December 31, 2021.

Account Title Debits Credits
Cash $ 40,000
Accounts receivable 170,000
Raw materials 44,000
Notes receivable 120,000
Interest receivable 23,000
Interest payable $ 25,000
Investment in debt securities 52,000
Land 70,000
Buildings 1,700,000
Accumulated depreciation—buildings 640,000
Work in process 62,000
Finished goods 109,000
Equipment 340,000
Accumulated depreciation—equipment 150,000
Patent (net) 140,000
Prepaid rent (for the next two years) 80,000
Deferred revenue 56,000
Accounts payable 200,000
Notes payable 600,000
Restricted cash (for payment of notes payable) 100,000
Allowance for uncollectible accounts 33,000
Sales revenue 1,200,000
Cost of goods sold 470,000
Rent expense 48,000


Additional Information:

  1. The notes receivable, along with any accrued interest, are due on November 22, 2022.
  2. The notes payable are due in 2025. Interest is payable annually.
  3. The investment in debt securities consist of treasury bills, all of which mature next year.
  4. Deferred revenue will be recognized as revenue equally over the next two years.


Required:
Determine the company’s working capital (current assets minus current liabilities) at December 31, 2021. (Amounts to be deducted should be indicated by a minus sign.)
Answer #1

Kansas Instruments corporation ($)

Working capital = Current assets - Current liabilities

= (Cash + Accounts receivable + Raw materials + Work in process + Finished goods + Notes receivable + Interest receivable + Short term investments + Prepaid rent) + (Interest payable + Accounts payable + Deferred revenue + Allowance for uncollected accounts)

= (40,000+170,000+44,000+62,000+109,000+120,000+23,000+52,000+40,000(80,000/2)) - (25,000+200,000+28,000(56,000/2)+33,000)

= 660,000 - 286,000 = $ 374,000

NOTES:

Investment in Debt securities are treated as short term investments, Current portion of prepaid rent and Deferred revenue are taken, Notes receivable and interest receivable are considered as current asset, Notes payable are long term in nature but interest is payable annually (current),
Answer #2
Allowance for uncollectible accounts is a negative current asset